An AM radio station broadcasts an electromagnetic wave with a frequency of 743 kHz, whereas an FM station broadcasts an electromagnetic wave with a frequency of 92.1 MHz. How many AM photons are needed to have a total energy equal to that of one FM photon?
1
Expert's answer
2020-10-12T07:50:29-0400

Since the energy of a photon is defined as "E=\nu\times h" , where "\nu" is the wave frequency and "h" is Planck's constant, to find the number of AM photons are needed to have a total energy equal to the energy of one FM photon, you need to find the ratio of energies of the FM to AM photon-photons:

"K=\frac { \nu_{FM}\times h }{ \nu_{AM}\times h }" "=\frac{ \nu_{FM}}{ \nu_{AM}}=" "\frac{ 92.1\times 10^6}{743\times 10^3}\approx123.96" , as the number of integer it is necessary "124" of the photon.